12x+7=9x+25
what is x? How do you figure it out

Answer by Theo(13342)

your equation is:

12x+7=9x+25

x represents the value you are looking for that will make the equation true.

this equation is true if the left side of the equation equals the right side of the equation.

that's because there is an equal sign between the left side of the equation and the right side of the equation.

you start with:

12x+7=9x+25

subtract 9x from both sides of the equation to get:

12x - 9x + 7 = 25

subtract 7 from both sides of the equation to get:

12x - 9x = 25 - 7

you have just moved all expressions with x in them to the left side of the equation and all constant expressions to the right side of the equation.

combine like terms to get:

3x = 18

divide both sides of the equation by 3 to get:

x = 18/3 = 6

the value of x that will satisfy the equation (make it true) is equal to 6.

substitute 6 for x in the original equation of 12x+7=9x+25 to get:

12*6 + 7 = 9*6 + 25

simplify by performing indicated operations to get:

72 + 7 = 54 + 25

combine like terms by performing indicated operations to get:

79 = 79

the left side of the equation is equal to the right side of the equation so the equation is true.

79 does equal 79.

if we set the value of x to 6, then the equation is true, so your answer is:

x = 6